# A Deep Learning-Based Supervised Transfer Learning Framework for DOA Estimation with Array Imperfections

### v2

@@ V-B2 Simulation results in Scenario 2 and 3
Figure 11: Scatter plots of our proposed model and MUSIC algorithm under Scenarios 4 with $snap=50$ and $SNR=-5db$. Scene 1 is configured with $\boldsymbol{\theta}=[-120,-60,0,60,120]$ and the corresponding azimuth angles $\boldsymbol{\phi}=[30,10,30,10,35]$. Scene 2 is configured with $\boldsymbol{\theta}=[-135,-120,-60,0,60]$, $\boldsymbol{\phi}=[35,30,10,30,10]$.
@@
Figure 12: The CDF of the absolute error under Scenarios 4 with $snap=50$ and $SNR=-5db$ is shown for: (a) $\boldsymbol{\theta}$, and (b) $\boldsymbol{\phi}$. The angles of the incident signals are randomly generated.
To evaluate the model’s performance in handling a larger number of incident angles or with more array elements, we extend the testing from Scenario 1 to Scenario 2 and Scenario 3.
We plotted the RMSE $vs.$ SNR results under Scenario 2 with $snap=10$ in Fig. 10. In this case, the incident signals are randomly generated within the range of $[-60^{\circ},60^{\circ}]$, and the minimum angular interval is set to 3 degrees. If the generated angular interval is less than 3 degrees, the incident angle will be regenerated. Notably, it can be observed that, compared with all the subspace-based methods, compressive sensing methods, and the state of the art deep learning algorithms, the proposed approach achieves superior statistical performance in low SNR and small snapshots scenarios. This indicates that the proposed model can still deliver relatively robust performance even in scenarios involving a greater number of incident angles.
Furthermore, the algorithm is tested under Scenario 3 with $\text{snap}=10$. In Fig. 10, it can be observed that the proposed model outperforms other algorithms in low SNR and small snapshot scenarios, except for cases where $SNR=-5db$ or $SNR=5db$. $\ell_{1}$-SVD and Unity-ESPRIT algorithms demonstrates high accuracy in scenarios with a larger number of array elements. Due to the limitations of data-driven algorithm fitting accuracy, the proposed model could not show significant improvement in algorithm performance in the range of 0 dB to 5 dB. As a result, its performance is inferior to Unity-ESPRIT and $\ell_{1}$-SVD algorithms under the $SNR=5db$ scenario. The performance of the proposed model in DOA estimation is difficult to improve in high-precision and higher SNR scenarios, which is an area for future improvement.
V-B3 Simulation results in Scenario 4
To evaluate the model’s ability to extend from one-dimensional to two-dimensional scenarios, we expand the model to two dimensions by simultaneously regressing the values of $\theta$ and $\phi$, and model’s performance is tested in Scenario 4.
Similar to one-dimensional DOA estimation, two-dimensional DOA estimation is simultaneously influenced by multiple parameters. Therefore, we consider two Scenes with different angular separations (AS3). Since it is uncommon to simultaneously vary the angular separations of all signals in practice scenarios, we only modify one of the angular separation parameters across the two scenes. In Scene 1, the incident signals are generated with $\boldsymbol{\theta}=[-120,-60,0,60,120]$, with the corresponding azimuth angles given by $\boldsymbol{\phi}=[30,10,30,10,35]$. Scene 2 is configured with $\boldsymbol{\theta}=[-135,-120,-60,0,60]$ and $\boldsymbol{\phi}=[35,30,10,30,10]$.
A representative visualization of the results is presented in Fig. 11, which facilitates an intuitive comparison of the algorithm’s performance. In Fig. 11, the symbol ”+” represents the true location of the target, while the dot in the figure represents the predicted value of the algorithm. Additionally, the model’s predicted values and their corresponding true values are marked with the same color. 100 Monte Carlo simulations were conducted for each scene. According to the experimental results in Scene 1, it can be observed that the predicted results of the proposed model are predominantly clustered around the true values. In contrast, the MUSIC algorithm achieves better DOA estimation performance at the more widely separated targets, like $\theta=[-120,120]$ and $\phi=[30,35]$, but it tends to confuse the other three points, resulting in poorer prediction performance. In Scene 2, we moved the fifth target from $\theta=[120]$ $\phi=[35]$ to $\theta=[-135]$ $\phi=[35]$, which is adjacent to the first target. It can be observed that the MUSIC algorithm fails to distinguish between the two targets, forming a single spectral peak between them, whereas the predicted results of the proposed model remain scattered around the true values.
@@ V-B3 Simulation results in Scenario 4
Figure 13: Three different levels of array imperfections are proposed to evaluate the performance of our proposed model and traditional algorithms: (a) $\rho=0$, no array imperfections, (b) $\rho=0.5$, moderate array imperfections, and (c) $\rho=1$, severe array imperfections.
Figure 14: RMSE $vs.$ SNR results under Scenarios 1 with $snap=10$ and array imperfections of $\rho=1$. The transfer learning approach is employed to enhance the model’s performance under array imperfection scenarios. The dashed line represents the case where only 50 samples are used for transferring learning, while the solid line indicates the use of 500 samples.
Furthermore, 1000 simulations are performed under Scenario 4 with $snap=50$ and $SNR=-5\ \mathrm{dB}$ to evaluate the statistical performance of the proposed model, where the angles of the incident signals are randomly generated. The CDF of the proposed model is presented in Fig. 12. It can be observed that the MUSIC algorithm performs better on CDF(10). However, its performance deteriorates on CDF(30). This suggests that while the MUSIC algorithm is likely to achieve more accurate DOA estimations for certain targets, it also shows a higher probability of producing significant DOA estimation errors. In contrast, the proposed model demonstrates a more consistent overall stability in its predictions.
This experiment also demonstrates that the CDF of Absolute Error serves as an effective metric in DOA estimation tasks. Compared with RMSE, it provides a more comprehensive evaluation of the model’s performance. Supported by the Cramér-Rao bound as a theoretical benchmark, RMSE offers a significant reference value, whereas the CDF of Absolute Error achieves a more thorough assessment of the model’s capabilities. Through the simulation results, it is evident that the proposed model achieves a greater advantage in terms of RMSE. However, through the analysis of the CDF of Absolute Error, it has certain limitations results in the precision DOA estimation.
@@ V-C Transfer Learning Approach for Array imperfection
The presence of array imperfections introduces inconsistencies between the received signal model and the ideal model, leading to a degradation in the DOA estimation performance of the model. Scenarios involving multiple array imperfections, particularly under low SNR and small snapshot conditions, are inherently complex and challenging to handle. In this section, we demonstrate the effectiveness of the proposed transfer learning approach in these scenarios through simulation experiments.
@@ V-C1 Transfer Learning in Scenario 1
Figure 15: DOA estimation under Scenarios 1 with $SNR=5db$ and $snap=10$, $\rho=1$. Three incident signals are generated and the interval is set to $\Delta\hat{\theta}={10}^{\circ}$. The line represents the true location of the target, while the dot in the figure represents the predicted value of the algorithm. Additionally, the model’s predicted values and their corresponding true values are marked with the same color.
TABLE II: Transfer Learning Results with Varying Sample Number
@@
(a) ViT without transfer learning
(b) ViT with transfer learning
Figure 16: DOA estimation under Scenarios 1 with $SNR=5db$, $snap=10$ and array imperfections of $\rho=1$. 100 Monte-Carlo simulations are conducted and the incident signal is configured with $\boldsymbol{\theta}=[-135,-120,-60,0,60]$, $\boldsymbol{\phi}=[35,30,10,30,10]$.
To demonstrate the performance degradation of the deep learning model under array imperfections, we illustrate the model’s performance under scenarios with multiple array imperfections, as depicted in Fig. 13. The position bias, amplitude and phase inconsistency, and mutual coupling effects of the array systems are simultaneously considered, and the intensity of these array imperfections is controlled by a hyperparameter $\rho$. As shown in Fig. 13, It can be observed that under the influence of multiple array imperfections, the statistical performance of deep learning-based methods (the proposed model, Learning-SPICE and SPE-CNN), compressed sensing-based methods ($\ell_{1}$-SVD), and subspace-based methods (Unity-ESPRIT and MUSIC) all deteriorates noticeably. Although the proposed model achieves optimal performance under the ideal signal model, particularly in low SNR and small snapshots scenarios, its performance significantly degrades when subjected to array imperfections. This degradation becomes particularly pronounced in scenarios with relatively higher SNR. When $\rho=1$, the performance of the proposed model declines level similar to the traditional algorithms at $\rho=0$.
To validate the effectiveness of the transfer learning approach, we evaluate the performance of the model adapted through transfer learning under varying SNR through 5,000 Monte Carlo simulations. As shown in Fig. 14, It can be observed that the proposed model achieves significant performance even with only 50 samples, demonstrating accurate DOA estimation under low SNR and small snapshot scenarios. Moreover, when more samples (500 samples) are used, the model’s performance is further enhanced.
To demonstrate the superiority of the proposed transfer learning algorithm, a comparison experiment is carried out under Scenario 1 with $snap=10$, $SNR=-5$, and array imperfections $\rho=1$. Our method is compared with two other approaches: (1) directly training the model through the samples with array imperfections (Directly Train), and (2) fine-tuning the model trained on simulated data using samples with array imperfections (Fine-tune). The performance of these three methods are evaluated across varying numbers of array imperfection samples, as shown in Table II.
@@
To illustrate the performance of the proposed model after transfer learning, an experiment is conducted to visually demonstrate our model’s performance, and the results are presented in Fig. 15. In this experiment, three incident angles are generated with a constant interval of $\Delta\hat{\theta}={10}^{\circ}$, covering the range from -60 to 60 degrees, with the incident angles linearly scaled with a step of 0.5 degrees. From the experimental results, it can be observed that the proposed model exhibits significant bias before applying transfer learning, particularly when the incident angles are close to -60 degrees. However, after applying transfer learning, the predict results of the proposed model become concentrated around the true values, demonstrating high DOA estimation accuracy. In contrast, due to the combined effects of multiple array imperfections, the performance of the MUSIC and $\ell_{1}$-SVD algorithm deteriorates, exhibiting substantial deviations. Meanwhile, Unity-ESPRIT and SPE-CNN show significant deviations when the incident angles span a wider range. SubspaceNet exhibits noticeable distortions, failing to achieve accurate DOA estimation. Learning-SPICE, SPE-CNN and ASL-2 demonstrate enhanced resilience to array imperfections; however, systematic distortion remains observable in the predict results.
@@ V-C2 Transfer Learning in Scenario 4
To further validate the effectiveness of the proposed transfer learning algorithm, we extend the algorithm from one-dimensional to two-dimensional. The 2D DOA estimation results of the ViT before and after applying transfer learning are shown in Fig. 16. It can be clearly observed that prior to the application of the transfer learning algorithm, the model fails to effectively distinguish between the first target ($\theta=[-135]$, $\phi=[35]$) and the second target ($\theta=[-120]$, $\phi=[30]$) and exhibits significant bias for orther targets due to array imperfections. However, after the transfer learning algorithm is applied, the prediction bias was substantially corrected, and the predicted values were able to converge near the true target.
@@ VI CONCLUSION
In this work, we develop a DOA-ViT model for DOA estimation, which leverages the attention mechanism to capture global dependencies. By directly learning the mapping from the sampled SCM to DOA values, the proposed method can be easily extended to arbitrary array geometries or 2D DOA estimation scenarios, demonstrating strong adaptability. Extensive simulations are conducted to evaluate the performance and robustness of the proposed model under various conditions: i) different SNRs and numbers of snapshots, ii) varying numbers of array elements $M$ and incident signals $K$, and iii) different angular separations and signal ranges.
Furthermore, deep learning based DOA estimators often suffer from severe performance degradation in practical scenarios with array imperfections. To overcome this limitation, we propose a supervised transfer learning strategy to mitigate the performance degradation in practical scenarios. By aligning the features of the source and target domains in a supervised manner, the proposed method demonstrates improved results in scenarios with array imperfections.
It is worth noticing that this work is the first to introduce an advanced transfer learning strategy to address array imperfections, rather than simply employing fine-tune. Investigating how to effectively leverage transfer learning methods to improve the performance of deep learning models in scenarios with array imperfections is a significant and worthwhile research direction, and has the potential to enhance the performance of deep learning algorithms in practical applications. Further exploration of various transfer learning approaches for handling array imperfections is encouraged.

### v5

@@ V-B2 Simulation results in Scenario 2 and 3
Figure 10: Scatter plots of our proposed model and MUSIC algorithm under Scenarios 4 with Deterministic configuration and Params [1].
To evaluate the model’s performance in handling increased source cardinality and larger array apertures, we conduct further experiments under Scenarios 2 and 3.
The RMSE $vs.$ SNR results under Scenario 2 is illustrated in Fig. 9. It can be observed that the proposed approach achieves superior statistical performance in low SNR and limited snapshots scenarios compared with all the subspace-based methods, compressive sensing techniques, and state of the art deep learning algorithms. This indicates that the proposed model maintains robustness even in high-density source environment.
Furthermore, the algorithm is evaluated under Scenario 3. As shown in Fig. 9, the proposed model outperforms other algorithms in low SNR scenarios, except for cases where SNR=5dbSNR=-5db or $SNR=5db$. Due to the limitations of data-driven algorithm fitting accuracy, the proposed model exhibits a performance plateau in the range of 0 dB to 5 dB. Consequently, its performance is inferior to Unity-ESPRIT and $\ell_{1}$-SVD algorithms under the $SNR=5db$ scenario, as these analytical methods benefit more effectively from the increased number of array elements. The performance of the proposed model in DOA estimation is difficult to improve in high-precision scenarios, which is an area for future improvement.
V-B3 Simulation results in Scenario 4
To evaluate the model’s ability in two-dimensional scenarios, the proposed architecture is extended to jointly regress the values of elevation $\boldsymbol{\theta}$ and azimuth $\boldsymbol{\phi}$ under Scenario 4.
Similar to one-dimensional DOA estimation, two-dimensional DOA estimation is inherently dependent on multiple factors. Therefore, we consider two DOA configurations with different angular separations. To facilitate a controlled comparison, we only modify one of the angular separation parameters between the two setups. Configuration 1 consists of incident signals at $\boldsymbol{\theta}=[-120,-60,0,60,120]$ with corresponding azimuth angles given by $\boldsymbol{\phi}=[30,10,30,10,35]$. In Configuration 2, the fifth target is relocated to test the resolution limits, resulting in $\boldsymbol{\theta}=[-135,-120,-60,0,60]$ and $\boldsymbol{\phi}=[35,30,10,30,10]$.
A representative visualization of the results is presented in Fig. 10, the symbol ”+” represents the true location of the target, while the dot in the figure represents the predicted value of the algorithm. Each predicted value shares the same color as its corresponding ground truth for clarity. Based on 100 Monte Carlo trials per configuration, it can be observed in Configuration 1 that the predicted results of the proposed model are predominantly clustered around the true DOAs. In contrast, while the MUSIC algorithm performs adequately for widely separated targets like $\theta=[-120,120]$ and $\phi=[30,35]$, but it tends to confuse the other three points, leading to degraded precision. In Configuration 2, the fifth target is moved adjacent to the first. It can be observed that the MUSIC algorithm fails to distinguish between the two targets, producing a single merged spectral peak. In contrast, the proposed model maintains high resolution, with predicted values still well clustered around their respective ground truth positions. These results highlight the superior capability of the model in achieving fine-grained localization in complex 2D DOA estimation tasks.
@@ V-B3 Simulation results in Scenario 4
Figure 11: The CDF of the absolute error under Scenarios 4 with $snap=50$ and SNR=5dbSNR=-5db is shown for: (a) $\boldsymbol{\theta}$, and (b) $\boldsymbol{\phi}$.
@@
Figure 12: RMSE $vs.$ SNR results under Scenarios 1 with Uniform configuration and Params [2]. Three different levels of array imperfections are proposed: (a) $\rho=0$, no array imperfections, (b) $\rho=0.5$, moderate array imperfections, and (c) $\rho=1$, severe array imperfections.
Figure 13: RMSE $vs.$ SNR results under Scenarios 1 with Uniform configuration, Params [2] and $\rho=1$. The dashed line represents the case where only 50 samples are used for transferring learning, while the solid line indicates the use of 500 samples.
Furthermore, 1000 simulations are performed under Scenario 4 with Uniform configuration and Params [1]. The ECDF of the absolute DOA estimation errors is presented in Fig. 11. It can be observed that the MUSIC algorithm performs better on $\text{CDF}(5^{\circ})$. However, its performance deteriorates on $\text{CDF}(15^{\circ})$. This suggests that although MUSIC can achieve high-precision estimations for certain samples, it is more prone to large estimation outliers. In contrast, the proposed model exhibits superior consistency and robustness.
This experiment further validates the effectiveness of ECDF as a comprehensive metric for DOA estimation. While RMSE remains a critical benchmark, especially when referenced against the CRLB, it only captures the average performance. In comparison, the ECDF provides a more detailed perspective on the error distribution. The results demonstrate that while the proposed model holds a clear advantage in terms of RMSE, the ECDF analysis reveals its subtle limitations in achieving high precision compared to MUSIC’s best-case scenarios.
@@ V-C Transfer Learning Approach for Array imperfection
The presence of array imperfections introduces inconsistencies between the received signal model and the ideal model, leading to a degradation in the DOA estimation performance of the model. Scenarios involving multiple array imperfections, particularly under low SNR and limited snapshot conditions, are inherently complex and challenging to handle. In this section, we demonstrate the effectiveness of the proposed transfer learning approach in these scenarios through simulation experiments.
@@ V-C1 Transfer Learning in Scenario 1
Figure 14: DOA estimation under Scenarios 1 with $SNR=5db$, $snap=10$ and $\rho=1$. The interval is set to $\Delta\hat{\theta}={10}^{\circ}$. The line represents the true location of the target, while the dot represents the predicted value of the algorithm.
TABLE III: Transfer Learning Results with Varying Sample Number
@@
(a) TransDOA without transfer learning
(b) TransDOA with transfer learning
Figure 15: DOA estimation under Scenarios 1 with Equidistant configuration, $SNR=5db$, $snap=10$ and $\rho=1$. 100 Monte-Carlo simulations are conducted and the incident signal is configured with $\boldsymbol{\theta}=[-135,-120,-60,0,60]$, $\boldsymbol{\phi}=[35,30,10,30,10]$.
To demonstrate the performance degradation of the deep learning model under array imperfections, we illustrate the model’s performance under scenarios with multiple array imperfections in Fig. 12. These imperfections include position bias, amplitude and phase inconsistency, and mutual coupling effects, with their overall severity controlled by a hyperparameter $\rho$. As shown in Fig. 12, It can be observed that under the influence of multiple array imperfections, the statistical performance of deep learning-based methods (the proposed model, Learning-SPICE and SPE-CNN), compressed sensing-based methods ($\ell_{1}$-SVD), and subspace-based methods (Unity-ESPRIT and MUSIC) all deteriorates noticeably. Although the proposed model achieves optimal performance under the ideal signal model, particularly in low SNR and limited snapshots scenarios, its performance significantly degrades when subjected to array imperfections. This degradation becomes particularly pronounced in scenarios with relatively higher SNR. When $\rho=1$, the performance of the proposed model declines level similar to the traditional algorithms at $\rho=0$.
The efficacy of the transfer learning scheme is further validated through 5,000 Monte Carlo trials across varying SNR levels. As shown in Fig. 13, the proposed model achieves significant performance even with only 50 samples, demonstrating accurate DOA estimation under low SNR and limited snapshot scenarios. Moreover, when more samples (500 samples) are used, the model’s performance is further enhanced.
To demonstrate the superiority of the proposed transfer learning algorithm, a comparison experiment is carried out under Scenario 1 with $snap=10$, SNR=5SNR=-5 (Params [1]), and array imperfections $\rho=1$. Our method is compared with two other approaches: (1) Direct Train, which involves training from scratch on imperfect data, and (2) Fine-tune, where a pre-trained model on simulated data is adapted samples with array imperfections. The performance of these three methods are evaluated across a varying number of training samples, as summarized in Table III.
@@
The impact of transfer learning is further highlighted in Fig. 14, considering the Equidistant configuration of Scenario 1 with $SNR=5db$, $snap=10$. In this experiment, three DOAs are generated with a constant interval of $\Delta\hat{\theta}={10}^{\circ}$, covering the range from -60 to 60 degrees. Evidently, the proposed model exhibits significant bias before applying transfer learning, particularly when the DOAs are close to -60 degrees. Transfer learning effectively rectifies these estimation errors, forcing the predicted estimates to converge tightly around the ground truth. In contrast, the performance of the MUSIC, $\ell_{1}$-SVD and SubspaceNet suffer from substantial deviations due to array imperfections. Meanwhile, Unity-ESPRIT and ASL-2 fail to maintain reliable estimation as the angular range expands. Learning-SPICE and SPE-CNNASL-2 demonstrate enhanced resilience to array imperfections; however, systematic distortion remains in the predict results.
@@ V-C2 Transfer Learning in Scenario 4
To further validate the scalability of the transfer learning approach, we extend its application from one-dimensional to two-dimensional. Fig. 15 presents the DOA estimation results of the TransDOA both before and after applying transfer learning. It can be clearly observed that prior to the application of the transfer learning algorithm, the model fails to effectively distinguish between the first two targets ($\theta=[-135]$, $\phi=[35]$ and $\theta=[-120]$, $\phi=[30]$) and suffers from significant bias for other targets due to array imperfections. However, after the transfer learning algorithm is applied, the prediction bias was substantially corrected, and the predicted values were able to converge near the true target.
@@ VI CONCLUSION
In this work, we develop a TransDOA model for DOA estimation, which leverages the attention mechanism to capture global dependencies. By directly learning the mapping from the sampled SCM to DOA values, the proposed method can be easily extended to arbitrary array geometries or 2D DOA estimation scenarios, demonstrating strong adaptability. Extensive simulations are conducted to evaluate the performance and robustness of the proposed model.
Furthermore, deep learning based DOA estimators often suffer from severe performance degradation in practical scenarios with array imperfections. To overcome this limitation, we propose a supervised transfer learning strategy to mitigate the performance degradation in practical scenarios. By aligning the features of the source and target domains in a supervised manner, the proposed method demonstrates improved results in practical scenarios with array imperfections.
It is worth noticing that this work is the first to introduce an advanced transfer learning strategy to address array imperfections. Investigating how to effectively leverage transfer learning methods to improve the performance of deep learning models in scenarios with array imperfections is a significant and worthwhile research direction, and has the potential to enhance the performance of deep learning algorithms in practical applications. Further exploration of various transfer learning approaches for handling array imperfections is encouraged.